A neurologist studying stimulant uptake by neurons adds 25 mL of 1M HNO3 to a solution containing 100 mL of 0.5M amphetamine, a molecule used to treat ADHD and narcolepsy. What is the pH of the solution after the titration of 25 degrees C? Kb of amphetamine = 6.31x10^-5
Answer #1

amphetamine is a weak base

so presence of acid will make conjugate acid

mmol of acid = MV = 25*1 = 25 mmol of acid

mmol of base = MV = 100*0.5 = 50 mmol of base

so... after reaction

25 mmol of acid forms 25 mmol of conjugate acid

50-25 = 25 mmol of base left

so

pOH = pKb + log(conjugate/base)

pOH = -log(6.31*10^-5) + log(25/25)

pOH = 4.1999

pH = 14-4.1999 = 9.8001